# Threads of Subtlety: Detecting Machine-Generated Texts Through Discourse Motifs

### v1

@@ Appendix D Calculation of motif features
Appendix E Preliminary analyses on discourse motifs
We begin our preliminary analyses by examining the relationship between the proposed discourse motifs and influential edges identified by the GAT explainer (§E.1). Subsequently, we calculate the difference distributions of motifs across five domains to emphasize their distinctive features (§E.2).
E.1 Correlation between Motifs and Influential Edges for Explanation
@@
E.2 Difference Distributions of Motifs across Domains
Figure 12: Difference in single-triad motif distribution of machine-generated and human-written texts for five different domains in DeepfakeTextDetect dataset. Y-axis represent the difference in MF-IDF scores scaled by 1e-3. Full-scaled versions including a larger number of motifs and plots for double- and triple-triads can be found in Appendix K.
@@ E.2 Difference Distributions of Motifs across Domains
Appendix F Constructing “TenPageStories” dataset
@@
Appendix G Details on experimental setup
@@
Appendix H More examples of grounded motifs in texts
@@
Appendix I Additional experiment: branched vs. chain-like structures
@@
Appendix J Difference distributions computed by motif frequency vs. MF-IDF scores
@@
Appendix K Large-scale difference distributions
@@
Appendix L Graphical examples of identified motifs

### v2

@@ Appendix C Calculation of motif features
Appendix D Preliminary analyses on discourse motifs
We begin our preliminary analyses by examining the relationship between the proposed discourse motifs and influential edges identified by the GAT explainer (§D.1). Subsequently, we calculate the difference distributions of motifs across five domains to emphasize their distinctive features (§D.2).
D.1 Correlation between Motifs and Influential Edges for Explanation
@@
D.2 Difference Distributions of Motifs across Domains
Figure 12: Difference in single-triad motif distribution of machine-generated and human-written texts for five different domains in DeepfakeTextDetect dataset. Y-axis represent the difference in MF-IDF scores scaled by 1e-3. Full-scaled versions including a larger number of motifs and plots for double- and triple-triads can be found in Appendix J.
@@ D.2 Difference Distributions of Motifs across Domains
Appendix E Constructing “TenPageStories” dataset
@@
Appendix F Details on experimental setup
@@
Appendix G More examples of grounded motifs in texts
@@
Appendix H Additional experiment: branched vs. chain-like structures
@@
Appendix I Difference distributions computed by motif frequency vs. MF-IDF scores
@@
Appendix J Large-scale difference distributions
@@
Appendix K Graphical examples of identified motifs